# MapAnything: Universal Feed-Forward Metric 3D Reconstruction map-anything.github.io

### v2

@@ Acknowledgments
0.14 0.03 86.7
0.14 0.02 86.3
0.01 0.02 94.2
0.01 0.01 92.9
Table S.2: MapAnything showcases versatile depth estimation under different input configurations on the Robust-MVD Benchmark [50]. Note that MapAnything has not been trained for single image inputs. In this full version of Table˜4, we additionally report the results for methods which produce up-to-scale or affine-invariant depth estimates, where they are aligned to the ground truth. We report the absolute relative error (rel) and the inlier ratio at a relative threshold of 1.03% ($\tau$). ‘K’ denotes camera intrinsics. The best result for each group is in bold; gray text indicates results where the evaluation dataset is in the training distribution [23].
KITTI ScanNet
a) Single-View Metric
MoGe-2 [70] 14.21 06.8 010.57 19.8
MapAnything 9.46 22.8 034.23 4.4
Depth Pro [5] 13.60 14.3 009.20 19.7
UniDepthV2 [44] 13.70 04.8 003.20 61.3
Metric3DV2 [19] 08.70 13.2 006.20 19.3
MapAnything 9.61 23.3 040.25 1.87
b) Multi-View Metric
MAST3R [30] 61.40 00.4 012.80 19.4
MUSt3R [6] 19.76 07.3 007.66 35.7
MapAnything 05.67 42.7 032.26 7.1
MapAnything 05.80 42.4 038.48 3.7
Fast-MVSNet [83] 12.10 37.4 287.10 09.4
MVS2D ScanNet [80] 73.40 00.0 004.50 54.1
MVS2D DTU [80] 93.30 00.0 051.50 01.6
Robust MVDB [50] 07.10 41.9 007.40 38.4
MAST3R Tri. [23] 03.40 66.6 004.50 63.0
MVSA [23] 03.20 68.8 003.70 62.9
MapAnything 04.23 56.4 017.31 11.9
c) Single-View w/ Alignment
MoGe [69] 05.12 46.2 003.59 65.3
MoGe-2 [70] 04.82 47.9 003.77 63.1
VGGT [66] 07.50 33.0 003.33 70.8
$\pi^{3}$ [74] 06.00 40.1 002.90 73.9
MapAnything 06.15 40.2 005.04 52.3
Depth Pro [5] 06.10 39.6 004.30 58.4
DAV2 [79] 06.60 38.6 004.00 58.6
Metric3DV2 [19] 05.10 44.1 002.40 78.3
UniDepthV2 [44] 04.00 55.3 002.10 82.6
MapAnything 06.09 41.0 004.70 55.2
d) Multi-View w/ Alignment
MAST3R [30] 03.30 67.7 004.30 64.0
MUSt3R [6] 04.47 56.7 003.22 69.2
VGGT [66] 04.60 53.0 002.34 80.6
$\pi^{3}$ [74] 03.09 69.5 001.98 83.6
MapAnything 04.07 58.0 003.96 60.7
DeMoN [60] 15.50 15.2 0012.00 21.0
DeepV2D KITTI [56] 03.10 74.9 023.70 11.1
DeepV2D ScanNet [56] 10.00 36.2 004.40 54.8
MapAnything 03.96 59.5 003.60 64.6
Figure S.1: The Apache and first stage (up to 4 views) training variants of MapAnything show strong dense multi-view reconstruction for number of input views varying from 2 to 100 and under different input configurations. We report the absolute relative error (rel), the inlier ratio at a relative threshold of $1.03\%$ ($\tau$), the average aligned trajectory error (ATERMSEATE~RMSE), the area under the curve at an error threshold of 5° ($AUC@5$), and the average angular error (errerr) in degrees (°), averaged over ETH3D, ScanNet++ v2 & TAv2.
Appendix A Implementation Details
[…]
Appendix B Additional Evaluation
Flexibility of Input Configurations:
We showcase a representative set of input configurations for MapAnything in Table˜S.1, where the performance of MapAnything improves as more modalities are provided. The universal training of MapAnything with varying selection probabilities for geometric inputs as 6 factors (as described in Section˜3.3) enables support for 64 exhaustive input combinations. While we primarily benchmark cases where the input modality is available for all views, MapAnything can also support optional geometric inputs for a subset of the input views.
Expanded Depth Benchmarking:
We provide an expanded version of Table˜4 in Table˜S.2, where the primary difference is the inclusion of results for methods which predict up-to-scale or affine-invariant depth estimates. Following the standard benchmark protocol, we report the alignment results for various methods using median alignment. We find that MapAnything provides versatile depth estimation performance across various settings where its performance is comparable to experts trained for specific tasks.
Comparison of MapAnything Variants:
We provide a comparison between different variants of MapAnything in Figure˜S.1, where the performance of VGGT [66] is provided as a baseline. Firstly, we showcase that our two-stage training with covisibility-based view sampling is very effective, where the MapAnything model trained for up to 4 views as input already shows strong generalization to a significantly higher number of views. We further also compare the performance of our different open-source models, where, as described in Table˜1, the Apache model is trained on 6 datasets, while the CC-BY-NC one is trained on 13. While we observe a decrease in performance, the Apache variant is still competitive to the VGGT baseline and its performance further improves as additional geometric inputs are provided.
@@ Appendix B Additional Evaluation
Design Choices:
As shown in Table˜S.3, log scaling and alternating attention are key for strong performance when evaluating with 50 views, well beyond the 4 used in training.

### v3

@@ Appendix C Additional Evaluation
0.12 0.02 92.0
0.12 0.02 83.0
0.01 0.02 94.9
0.01 0.02 84.4
@@
Figure S.5: Auxiliary geometric inputs improve feed-forward performance of MapAnything. (Top) While MapAnything & other baselines using 100 input images show duplication of 3D structure, when provided with the camera calibration and poses, the 3D reconstruction significantly improves, showing aligned geometry. (Middle) MapAnything using images only as input shows non-precise metric scale estimation on ETH3D (a zero-shot dataset). However, when the calibration and metric poses are provided as additional input, the estimated metric scale significantly improves and approximately matches the ground truth. (Bottom-Left) We show that MapAnything can leverage a sparse metric point cloud as input to perform dense metric depth completion. (Bottom-Right) Despite not being trained for object-centric data, we show how the scene geometry and cameras change based on the input provided.
Figure S.6: MapAnything showcases state-of-the-art and competitive dense multi-view reconstruction performance in comparison to latest public concurrent state-of-the-art models as of January 20th 2026, i.e., DA3-Nested & $\pi^{3}$-X, despite being trained on significantly less amount of data. We report the absolute relative error (rel), the inlier ratio at a relative threshold of $1.03\%$ ($\tau$), the average aligned trajectory error (ATERMSEATE~RMSE), the area under the curve at an error threshold of 5° ($AUC@5$), and the average angular error (errerr) in degrees (°), averaged over ETH3D, ScanNet++ v2 & TAv2.
Figure S.7: Dense multi-view reconstruction benchmarking across individual datasets using only images as input. We also include results for latest public concurrent state-of-the-art models as of January 20th 2026. Please see Figure˜S.6 for details & the averaged version.
Figure S.8: Dense multi-view reconstruction benchmarking across individual datasets using multi-modal inputs. We also include results for latest public concurrent state-of-the-art models as of January 20th 2026. Please see Figure˜S.6 for details & the averaged version.